HUBBLE SPACE TELESCOPE DAILY REPORT #1954 PERIOD COVERED: 0000Z (UTC) 09/02/97 - 0000Z (UTC) 09/03/97 Daily Status Report as of 246/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7659 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make dark calibration images in order to update the acquisition bad pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Three Sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.4 Completed Three Sets of STIS/CCD/MA1 7722 (Lamp Flux Measurements) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make lamp flux calibration measurements to determine the optimal configuration for obtaining science-contemporaneous wavecals. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed NIC/2 7330 (The Fueling of Active Nucleii: A NICMOS Snapshot Survey) The Near Infrared Camera (Camera Two) was used to probe the dark matter distribution of NGC-5347 far from the center by observing the intervening absorption from the MgII doublet (lambda 2800) in the spectrum of one background QSO located in a geometrically optimal position. HST observations are required because this feature is in a wavelength range not accessible from the ground, and the QSO magnitudes and expected strength of absorption make IUE observations unfeasible. The proposal completed as expected with no anomalies. 1.6 Completed NIC/1/2/3 7412 (An HST/NICMOS Search for Young Brown Dwarfs and Giant Planets) The Near Infrared Camera (Cameras One, Two, and Three) was used to make observations of RXJ12319-7 taken in parallel with the prime science observations to be part of a survey to look for large planets and/or small brown dwarf stars which are in orbit around T-Tauri class stars. At an age of 2 to 20 million years, these low-mass companions will be considerably brighter than more evolved brown dwarfs or giant planets orbiting main-sequence stars, and hence will be much easier to detect. The observations were completed as planned, and no anomalies were noted. 1.7 Completed WF/PC-2 7786 (Pure Parallel Archive Survey - Continuation) The WF/PC-2 was used to make observations as part of an archival survey program in which the images were taken in parallel with the prime science observations. These WF/PC-2 parallel observations were obtained in the ultraviolet in a search for interesting objects. The observations were completed as planned, and no problems were reported. 1.8 Completed WF/PC-2 6587 (Black Holes and Cores of Early-Type Galaxies) The WF/PC-2 was used to make observations of early-type galaxy NGC-596 in order to study its nuclear core region and its possible black hole. The campaign to find black holes in the centers of galaxies using HST has so far been highly successful. The observations were completed as planned, and no problems were reported. 1.9 Completed STIS/CCD 7641 (CCD External Flats, Stellar) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of NGC-104 (a globular cluster in Tucana) as part of an external stellar flat calibration observations. The purpose of this proposal is to measure accuracy of the CCD flat fielding around dust features on the filters. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed NIC/1/2/3 7447 (The Cosmological Evolution of Quasar Host Galaxies) The Near Infrared Camera (Cameras One, Two, and Three) was used to make observations of galaxy 1148+56W1 as part of a survey to study the cosmological evolution of galaxies which host quasars. Other point spread function observations were taken to test the calibration of the cameras. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed STIS/CCD/MA1/MA2 7667 (MAMA Geometric Distortion) The Space Telescope Imaging Spectrograph (CCD, and both MA 1 and 2) was used to make observations of NGC-346 (a cluster associated with a nebulosity in Tucana) in order to study the MAMA geometric distortion and derive the geometric distortion of MAMA first order spectra for those cases that were missed in ground calibrations. The technique is to step three bright stars along the wide slit, obtaining a high signal to noise spectrum at each displacement. Thirty points will be sampled within +/- 12" of the center. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Operations Notes: The STIS MCE1 reset outside of the South Atlantic Anomaly at 246/064912Z. The STIS MAMAs went from high voltage on to low voltage on, and occurred while the STIS 7659 proposal was executing. It is strongly suspected that this is a result of solar storm activity reported by SOHO. No action will occur until day shift when the recoevery will be planned. The next high voltage on is scheduled for 246/214858Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH